Vertify the identity
cosx(cosx+cotx)=(cotx.cosx)(sinx+1)

Answer by Alan3354(69443)

cosx(cosx+cotx)=(cotx.cosx)(sinx+1)
cos^2 + sin = (cos^2/sin)(sin + 1) = cos^2 + cos^2/sin
sin = cos^2/sin
sin^2 = cos^2
Not an identity.
Maybe cotx.cosx doesn't mean multiply?